### v2

# Federated Nonlinear System Identification

Abstract
We consider federated learning of linearly-parameterized nonlinear systems. We establish theoretical guarantees on the effectiveness of federated nonlinear system identification compared to centralized approaches, demonstrating that the convergence rate improves as the number of clients increases. Although the convergence rates in the linear and nonlinear cases differ only by a constant, this constant depends on the feature map $\phi$, which can be carefully chosen in the nonlinear setting to increase excitation and improve performance. We experimentally validate our theory in physical settings where client devices are driven by i.i.d. control inputs and control policies exhibiting i.i.d. random perturbations, ensuring non-active exploration. Experiments use trajectories from nonlinear dynamical systems characterized by real-analytic feature functions, including polynomial and trigonometric components, representative of physical systems including pendulum and quadrotor dynamics. We analyze the convergence behavior of the proposed method under varying noise levels and data distributions. Results show that federated learning consistently improves convergence of any individual client as the number of participating clients increases. footnotetext: * denotes equal contribution
1 Introduction
Dynamical system models describe how a system evolves over time based on its current state, control inputs, and external disturbances. These models are central to fields including control theory, physics, and robotics (Willems, 1989). A common class of such systems is linear time-invariant (LTI) systems, where the next state and output are linear functions of the current state and input, governed by parameters that remain constant over time (Sibai & Mitra, 2018). In contrast, nonlinear systems allow these dependencies to be governed by nonlinear functions, offering more accurate representations for many real-world systems with complex or nonstationary behavior (Guckenheimer & Holmes, 2013; Jiang & Lai, 2019).
Figure 1: Federated learning framework for nonlinear dynamical system identification involving $M$ clients which are similar but non-identical in nature and a central server. In each global communication round $r$, client $C_{i}$ receives the global model $\theta_{s}$ performs local updates using its own trajectories data, and transmits the locally updated model $\theta_{i}$ back to the server. The server then aggregates the models to obtain an updated global model for the next round.
System identification is the process of learning dynamical models from input and output data. In this way, one can learn the dynamics of a system in a data-driven way, without explicit physical derivations. It is often used to learn dynamical systems in the real world, where exact dynamics are unknown, with applications including robotics, fluid dynamics, and physics (Sarkar et al., 2021; Venkatesh & Dahleh, 2002). Traditional approaches for system identification typically assume access to centralized datasets. However, in modern deployments—such as multi-robot systems, sensor networks, and connected vehicles—data is naturally distributed across many devices. Transmitting all data to a central location is often impractical due to privacy, bandwidth, or energy limitations.
To address these challenges, we consider federated system identification, which is emerging as a privacy-preserving alternative. In this paradigm, multiple clients collaboratively learn a model without sharing raw data, communicating only model updates. Prior work has explored federated identification for LTI systems (Wang et al., 2023), demonstrating convergence benefits that scale with the number of participating clients. However, this line of work is restricted to linear models, which fundamentally limits its applicability.
In this paper, we study the problem of federated nonlinear system identification. We consider a setting where $M$ distributed clients each observe trajectory data from potentially different nonlinear dynamical systems that nonetheless belong to the same underlying family (see Figure 1). Dynamical systems at clients exhibit heterogeneity, modeled via a heterogeneity parameter $\epsilon$. This setup captures realistic scenarios such as learning dynamics across a fleet of heterogeneous robots or drones—each operating under different conditions due to hardware differences, controller tunings, or external disturbances. The goal is to identify a shared nonlinear model that generalizes across these clients while respecting local variability and preserving data privacy.
1.1 Key Contributions
Figure 2: Impact of client count and heterogeneity on federated nonlinear system identification. (a) Estimation error versus number of clients $M$ at fixed trajectories per client $N_{i}=10$ and heterogeneity bound $\epsilon=0.01$; (b) Estimation error versus $M$ for the same $N_{i}$ with $\epsilon=5$, illustrating increased model-mismatch tolerance; (c) Estimation error as a function of $\epsilon$ for fixed $M$ and $N_{i}$, showing sensitivity of global model accuracy to client-level parameter variations.
  • Nonlinear framework. We consider linearly parameterized nonlinear dynamical systems, with a focus on the widely used piecewise affine (PWA) model for representing nonlinear dynamics (Mania et al., 2022; Musavi et al., 2024). In PWA models, state transitions are expressed as linear functions of nonlinear feature embeddings of state-input pairs. While prior work has explored federated system identification of linear dynamical systems, to the best of our knowledge, this work develops the first federated learning framework for system identification of PWA-based nonlinear dynamical systems.
  • Theoretical analysis. We theoretically analyze the convergence of the federated nonlinear system identification problem. Our results show that clients benefit significantly by collaborating through federated learning. Particularly, we show that the convergence error decreases as $\mathcal{O}(1/\sqrt{M})$, where $M$ is the number of clients, thus having significant improvement in convergence as more clients collaborate. Additionally, our results theoretically characterize how heterogeneity across clients impacts the convergence error.
  • Experimental analysis. To corroborate our theoretical results, we conduct extensive experiments analyzing the convergence behavior of our approach as a function of the total number of participating clients, the number of local samples at each client, as well the heterogeneity of dynamical systems across clients. We preview some results in Figure 2, which are based on real-world pendulum dynamics. Complete details are provided in Section 4.
  • Algorithm validation. Our experiments demonstrate that our federated approach achieves improved convergence rate compared to a single-client system, while offering the additional advantage of client data privacy.
2 Related Works
2.1 System Identification
Ljung (1998) provides an overview of mathematical methods for building models of dynamical systems using input and output signal measurements, focusing on how parameter estimation techniques can be applied to characterize system behavior in both time and frequency domains. This approach is instrumental in transforming observed data into predictive models, supporting applications in signal processing and control. Jedra & Proutiere (2020) present a robust finite-time error bound for identifying stable linear systems using the ordinary least-squares (OLS) estimator. Simchowitz et al. (2018) also analyze OLS identification from a single trajectory for stable and marginally stable linear time-invariant systems, using the block-martingale small-ball method to address mixing dependencies and establish nearly optimal estimator rates for a wide class of linear time series. While deriving tight, regime-specific finite-time error rates for OLS identification of general linear systems, Sarkar & Rakhlin (2019) establish that OLS is optimally consistent only under specific regularity conditions, unifying previous results across all spectral regimes. However, this approach struggles with nonlinear or complex signal structures, often requiring significant computational resources and large datasets to achieve reliable identification.
2.2 Federated Learning
Federated learning is a machine learning paradigm which allows clients to collaboratively train a model without sharing raw data (McMahan et al., 2017). In principle, each client possesses a local dataset and sends weight updates to the server. The server then aggregates these weights together to update the global model. This approach is very useful to train models on sensitive or private data. Truex et al. (2019) propose a hybrid federated learning approach that combines differential privacy and secure multiparty computation to enhance privacy protection while maintaining high model accuracy in collaborative machine learning settings. The paper presents communication-efficient federated learning strategies through structured low-dimensional updates and compressed sketched updates, achieving a substantial reduction in communication overhead while preserving model accuracy (Konečnỳ et al., 2016). Li et al. (2020) introduce the FedProx framework, which adds a proximal term to FedAvg and provides convergence guarantees to achieve more robust and stable training in the presence of statistical and system heterogeneity across clients. Federated learning architectures can use a peer-to-peer approach (Yang et al., 2019). However, our discussion centers on the traditional server-client model, in which a central server periodically aggregates updates from multiple clients. While most federated learning works are concerned with learning deep neural networks, our work focuses on understanding the theory and implementation of federated system identification of nonlinear dynamical systems.
2.3 Federated System Identification
The goal of the federated system identification problem is to determine a shared estimate of the system matrices, denoted as $\bar{\theta}$, that yields a small estimation error relative to each client $i$’s true system parameters $\theta^{(i)*}$. Clients $i$ do not share their private data with the server, and each client independently estimates its unknown system dynamics. Wang et al. (2023) formulated an FL framework for system identification of linear system models across heterogeneous clients with improved sample efficiency. They show that more clients improve the model convergence rate and propose the FedSysID algorithm. Their work was restricted to synthetic linear systems, which limits applicability to real-world nonlinear dynamical systems. This limitation motivates our proposal of federated nonlinear system identification, as detailed in the following sections.
3 Federated Nonlinear System Identification
In this section, we formally define the federated nonlinear system identification framework. Then we provide a convergence rate of the system with respect to the number of clients, number of trajectories at each client, trajectory length, and heterogeneity across clients. Finally, we present the FNSysId algorithm that describes the client-server interactions for nonlinear system identification.
3.1 Preliminaries
The formal objective of the federated system identification problem is to estimate a model that accurately captures the collective behavior of the participating systems. Specifically, suppose that for each client $i\in[M]$, we have the following linearly parameterized nonlinear dynamics:
$x_{t+1}^{(i)}=\theta^{(i)*}\phi(x_{t}^{(i)},u_{t}^{(i)})+w_{t}^{(i)},$ (1)
where $x_{t}^{(i)}$, $u_{t}^{(i)}$, and $w_{t}^{(i)}$ denote the state, control, and disturbance respectively. Furthermore, $\phi:\mathbb{R}^{n_{x}}\times\mathbb{R}^{n_{n}}\to\mathbb{R}^{n_{\phi}}$ denotes a vector of known nonlinear mappings (Mania et al., 2022; Musavi et al., 2024).
In the following, we recall two key definitions from Musavi et al. (2024) that will be used in our assumptions.
Definition 1.
(Semi-continuous distribution). A probability distribution $\mathbb{P}$ is semi-continuous if there does not exist a set $E$ with Lebesgue measure zero such that $\mathbb{P}(E)=1$.
Definition 2.
[…]
We next provide four assumptions for the local dynamical systems at clients (Musavi et al., 2024) (we omit client indexing for brevity):
  • Assumption 1: All components of the feature vector $\phi(\cdot)$ are real-analytic functions in $\mathbb{R}^{n_{x}+n_{u}}$, i.e., all components in $\phi(\cdot)$ are infinitely differentiable.
  • Assumption 2: Noise $w_{t}$ is i.i.d., and follows a semi-continuous, mean-zero distribution with positive semi-definite covariance matrix $\sum_{w}\succeq\sigma_{w}^{2}I_{n_{x}}$.
  • Assumption 3: Input $u_{t}$ is i.i.d following a semi-continuous, mean-zero distribution with positive semi-definite covariance matrix $\sum_{u}\succeq\sigma_{u}^{2}I_{n_{x}}$.
[…]
Additionally, inspired by Wang et al. (2023), we consider the following assumption of bounded system heterogeneity across clients.
[…]
Our goal is to solve the least squares error estimate problem of our system to find the optimal system parameters:
$\bar{\theta}_{LSE}=\frac{1}{M}\sum_{i=1}^{M}\arg\min_{\theta\in\mathbb{R}^{n_{\phi}\times n_{x}}}\bigl{\|}X_{+}^{(i)}-\theta\,\Phi^{(i)}\bigr{\|}_{F}^{2},$ (2)
where $X_{+}^{(i)}=\bigl{[}x_{0}^{(i,1)},\dots,x_{T-1}^{(i,N_{i})}\bigr{]}$, and $\Phi^{(i)}\in\mathbb{R}^{n_{\phi}\times(N_{i}T)}$ is defined as follows:
$\Phi^{(i)}=\bigl{[}\phi(x^{(i,1)}_{0},u^{(i,1)}_{0}),\dots,\phi(x^{(i,N_{i})}_{T-1},u^{(i,N_{i})}_{T-1})$].
Functionally, this optimization estimates each client’s local dynamics via least-squares regression, then averages the results to obtain a global model.
3.2 Framework Convergence Rate
A least-squares estimate is meaningful only if the design matrix built from each client’s regressor vectors is well conditioned. In linear system identification this role is played by a persistent-excitation assumption. For linearly parameterized nonlinear dynamics, we recover a similar property through the block-martingale small-ball (BMSB) condition. Lemma 1 is used to show that the analytic feature map and new inputs guarantee the required excitation.
Lemma 1 (BMSB for open-loop systems).
Let each client $i\in[M]$ run open-loop inputs $u^{(i)}_{t}=\eta^{(i)}_{t}$. Let client $i$ collect a trajectory of length $T$. For every $t=0,1,\dots,T-1$, define the filtration
$\mathcal{F}^{(i)}_{t}:=\sigma\bigl{(}w^{(i)}_{0},\dots,w^{(i)}_{t-1},\;x^{(i)}_{0},\dots,x^{(i)}_{t},\;\eta^{(i)}_{0},\dots,\eta^{(i)}_{t}\bigr{)}$.
Under Assumptions 1–4 there exist constants $s_{\phi}>0$ and $p_{\phi}\in(0,1)$ (as defined by Musavi et al. (2024)) such that, for every client $i$ and every unit vector $v\in\mathbb{S}^{n_{\phi}-1}$,
$\Pr\!\bigl{(}|v^{\!\top}\phi(x^{(i)}_{t},u^{(i)}_{t})|\,\geq s_{\phi}\;\big{|}\;\mathcal{F}^{(i)}_{t-1}\bigr{)}\;\geq\;p_{\phi}\quad\text{a.s.}$ (3)
Hence the regressor process $\{\phi(x^{(i)}_{t},u^{(i)}_{t})\}_{t\geq 0}$ satisfies the $(1,s_{\phi}^{2}I_{n_{\phi}},p_{\phi})$-BMSB condition for every client $i$.
Proof.
Please refer to the Appendix for the proof. ∎
Less formally, each new regressor $\phi(x^{(i)}_{t},u^{(i)}_{t})$ has (conditioned on the past) at least probability $p_{\phi}$ of having length $s_{\phi}$ in every direction, hence the excitation we desired.
To use least-squares, the empirical Gram matrices must be invertible. Proposition 1 provides that guarantee and will be used to determine the $1/N_{\mathrm{tot}}$ factor in our final error bound.
Proposition 1.
Fix $\delta\in(0,1)$. Suppose that each client $i$ collects $N_{i}$ samples of length $T$. For every client $i\in[M]$, where $N_{\mathrm{tot}}:=\sum_{i=1}^{M}N_{i}T$, if each sample size satisfies $N_{i}T\;\geq\;\frac{4}{p_{\phi}}[n_{\phi}\log(9)+\log(M/\delta)],$ then, with probability at least $1-\delta$,
$\lambda_{\min}\!\bigl{(}\Phi^{(i)}\Phi^{(i)\!\top}\bigr{)}\;\geq\;\tfrac{1}{2}\,s_{\phi}^{2}N_{i}T,\quad i\in[M],$
and hence the pooled Gram matrix
$G:=\sum_{i=1}^{M}\Phi^{(i)}\Phi^{(i)\!\top}\succeq\tfrac{1}{2}\,s_{\phi}^{2}N_{\mathrm{tot}}\,I_{n_{\phi}}.$ (4)
Proof.
Please refer to the Appendix for the proof. ∎
We now need to bound the stochastic noise regressor cross-term. This will be used to guarantee that the estimator remains stable and accurate despite disturbances.
Proposition 2.
Fix $0<\delta<1$. For client $i\in[M]$ define the noise matrix
$W^{(i)}:=[w^{(i)}_{0}\;w^{(i)}_{1}\;\dots\;w^{(i)}_{N_{i}-1}]\in\mathbb{R}^{n_{x}\times N_{i}T},$
and design matrix
$\Phi^{(i)}:=\big{[}\phi(x^{(i,j)}_{t},u^{(i,j)}_{t})\big{]}_{\begin{subarray}{c}j=1,\dots,N_{i}\\[2.0pt] t=0,\dots,T-1\end{subarray}}\in\mathbb{R}^{n_{\phi}\times N_{i}T}.$
Then, with probability at least $1-\delta$,
$\|W^{(i)}(\Phi^{(i)})^{\top}\|_{2}\leq 4\sigma_{w}\sqrt{\,N_{i}T(n_{x}+n_{\phi}+\log(2M/\delta))}$ (5)
for $i\in[M]$.
Consequently, for the pooled noise matrix $P:=\sum_{i=1}^{M}W^{(i)}(\Phi^{(i)})^{\!\top},$ the same event implies
$\|P\|_{2}\;\leq\;4\sigma_{w}\sqrt{\,N_{i}T(n_{x}+n_{\phi}+\log(2M/\delta))}.$ (6)
Proof.
Please refer to the Appendix for the proof. ∎
These auxiliary propositions can now be combined to bound the final convergence rate.
Theorem 1 (Finite-sample error).
Let $s_{\phi}$, $p_{\phi}$ be as in Lemma 1 and set
$C_{1}:=\frac{8\sigma_{w}}{s_{\phi}^{2}},\qquad\,C_{2}:=\frac{b_{\phi}}{s_{\phi}^{2}}+\frac{1}{2}.$
Then for any $\delta\in(0,1)$, with probability at least $1-3\delta$,
$\bigl{\|}\bar{\theta}_{\mathrm{LSE}}-\theta^{(i)\!*}\bigr{\|}_{2}\;\leq\;C_{1}\sqrt{\frac{n_{x}+n_{\phi}+\log(2M/\delta)}{T\sum_{i=1}^{M}N_{i}}}+C_{2}\,\varepsilon$
for all $i\in[M].$
Proof.
Please refer to the Appendix for the proof. ∎
Remark 1.
The first term in the convergence error decreases as $1/\sqrt{T\sum_{i=1}^{M}N_{i}}$, where $\sum_{i=1}^{M}N_{i}$ is the total number of trajectories across all clients. Furthermore, since the number of trajectories $N_{i}$ at client $i$ is a constant for a given trajectory length $T$, the error decreases as $\mathcal{O}(1/\sqrt{M})$.
Remark 2.
For smaller $\epsilon$, the underlying system dynamics at the clients are more similar, enabling FL clients to effectively use data from all participants. This enhances the performance of each client, particularly achieving a convergence speedup of $\sqrt{M}$.
3.3 FNSysId
The core workflow of FNSysId implementation is formalized in Algorithm 1, which shows how clients perform local computations and synchronize model updates with the central server during each communication round.
Algorithm 1 FNSysId
Initialize the server with $\bar{\theta}_{0}$, and learning rate $\alpha$ ;
Initialize each client $i\in[M]$ with $\theta_{0,0}^{{(i)}}=\bar{\theta}_{0}$;
For each global iteration $r=0,1,\ldots,R-1$ do
[…]
  For each client $i\in[M]$ in parallel do
   $\theta_{r+1}^{{(i)}}=\texttt{ClientUpdate}(i,\bar{\theta}_{r},K_{i})$
   send $\theta_{r+1}^{{(i)}}$ back to the server
  end for
[…]
  update $\bar{\theta}_{r+1}=\frac{1}{M}\sum_{i=1}^{M}\theta_{r+1}^{{(i)}}$ and send $\bar{\theta}_{r+1}$ to each client
end for
Return $\bar{\theta}_{R}$
Each client initializes a guess for its local model parameters, denoted as $\bar{\theta}_{0}$, and a learning rate $\alpha$. These local initializations are critical for algorithm convergence and allow each client to tailor its optimization process based on its individual system characteristics and data availability.
In our framework, we assume all clients $\mathcal{C}=\{C_{1},C_{2},\ldots,C_{M}\}$ actively participate in every global iteration $r\in[R]$. Each client performs $K_{i}$ local updates independently using its own local data before any communication with the central server (line 6), where ClientUpdate iteratively executes:
$\theta_{r,k}^{{(i)}}=\;$ $\theta_{r,k-1}^{{(i)}}+\alpha\left({y}^{{(i)}}-\theta_{r,k-1}^{{(i)}}{X}^{{(i)}}\right)({X}^{{(i)}})^{\top},\quad k=1,2,\ldots,K_{i}.$ (7)
This approach reduces communication overhead by allowing multiple updates locally. After completing these updates, the clients send their updated local models $\bar{\theta}_{r+1}^{(i)}$ to the server. The server then averages all updates to obtain a new global model $\bar{\theta}_{r+1}$, which is subsequently broadcast back to all clients to synchronize (line 10). This efficiently balances local computation and communication, addressing potentially limited communication capabilities between clients.
4 Experiments
We conducted experiments on a range of systems, encompassing both synthetic benchmarks and real-world dynamical systems. In the following, we provide details regarding datasets employed, hyperparameter configurations, and the corresponding results.
4.1 Data
Conducting experiments on both synthetic and real-world system data is essential to comprehensively evaluate the proposed methodology. Synthetic datasets offer controlled environments enabling precise assessment of model accuracy and convergence behavior. In contrast, real-world datasets provide insights into the method’s practical applicability, robustness to noise, and generalization capabilities under complex, uncontrolled conditions. This dual evaluation ensures both theoretical soundness and empirical relevance.
4.1.1 Synthetic Data
To evaluate the performance and efficiency of FNSysId, we conduct numerical experiments involving $M$ clients, each associated with a discrete-time dynamical system characterized by $n=3$ states and $p=2$ inputs. The initial state, input signal, and process noise are all sampled from zero-mean distributions with unit standard deviations. The trajectories were generated by augmenting the trajectory formulation presented by Wang et al. (2023) with a nonlinear term, as presented in the following equation.
$x_{t+1}^{(i)}=A^{(i)}\sin(x_{t}^{(i)})+B^{(i)}u_{t}^{(i)}+w_{t}^{(i)}.$ (8)
Following the approach of Xin et al. (2022), we construct heterogeneous client dynamics by perturbing a nominal system $(A_{0},B_{0})$. Particularly, for client $i$, the perturbed system matrices are as follows:
$A^{(i)}=A_{0}+\gamma_{1}^{(i)}V,$ (9)
$B^{(i)}=B_{0}+\gamma_{2}^{(i)}U,$ (10)
where $\gamma_{1}^{(i)}\sim\mathrm{U}(0,\epsilon)$ and $\gamma_{2}^{(i)}\sim\mathrm{U}(0,\epsilon)$ are uniformly distributed random variables.
We consider a trajectory length of $T=5$ and assume that each client uses an equal number of data points.
4.1.2 Real-World System
We next conduct experiments across two different physical systems whose characteristics and control requirements provide contrasting evaluation scenarios. The first is a pendulum system characterized by one state and one input, targeting the estimation of two unknown parameters. Further details on the trajectory generation methodology are provided in Section 4 of Musavi et al. (2024). In all experimental settings, the control input $u_{t}$ is defined as the superposition of a deterministic policy and stochastic noise, i.e., $u_{t}=\pi(x_{t})+\eta_{t},$ where $\pi(x_{t})$ denotes the control policy evaluated at state $x_{t}$ and $\eta_{t}$ represents zero-mean i.i.d. noise. Here $\eta_{t}$ is sampled from uniform distribution distribution to model bounded stochastic disturbances.
For the pendulum, we have two nominal parameters: $A_{0}=\tfrac{1}{l}$ and $B_{0}=\tfrac{1}{ml^{2}}$. Here, $m$ denotes the mass of the pendulum and $l$ represents its length. Following the formulation of Musavi et al. (2024), we modified the trajectories by introducing the heterogeneity using the same procedure outlined in equation 9 and equation 10.
$\ddot{\alpha}=-A^{(i)}g\sin(\alpha)+B^{(i)}u+w,$ (11)
where $\alpha$ represents the angle of the rod relative to the vertical axis, $g$ is the gravity constant, $u$ is the torque input, and $w$ is the disturbance applied to this system. After discretization, the system dynamics can be rewritten in the structure of (1) with the feature vector consisting of expressions involving $\sin(\alpha)$ and $u$, all of which are analytic functions. The matrix of unknown parameters contains terms of the pendulum’s mass and the rod’s length. Each trajectory consists of $T=5$ time steps.
In the quadrotor example presented in Section 4 of Musavi et al. (2024), the control input is specified as $u_{t}=\pi(x_{t})+\eta_{t}$, where $\pi(x_{t})$ denotes the nominal controller designed according to the approach in Alaimo et al. (2013). The system comprises $13$ states and $4$ control inputs. The unknown parameter matrix $\theta_{*}$ includes $7$ parameters, such as the mass $m$ and selected elements of the inertia matrix $I$.
4.2 Estimation Error
For each client $C_{i}$, the normalized error between the client and the server parameters is calculated as:
$e_{r}^{(i)}=\frac{\left\|\theta_{s,r}-\theta^{(i)*}\right\|_{2}}{\left\|\theta^{(i)*}\right\|_{2}}.$ (12)
The maximum estimation error reported for communication round $r$ is: $e_{r}=\max_{1\leq i\leq M}e_{r}^{(i)}$, which also serves as a heterogeneity metric that quantifies how far the global model is from the most divergent client.
5 Results
Figure 3: Estimation error versus the number of global iterations for the real-world nonlinear dynamical system of a pendulum using gradient descent(GD). Results illustrate the impact of varying the number of clients ($M$), the number of local samples per client ($N_{i}$), and the heterogeneity parameter ($\epsilon$), with each client performing $K_{i}=1$ local updates at alearning rate of $10^{-2}$. Subfigures: (a) $N_{i}=10$, $\epsilon=0.01$; (b) $M=10$, $\epsilon=0.01$; (c) $M=20$, $N_{i}=10$.
Figure 4: Estimation error vs. global iterations for the nonlinear pendulum system using mini-batch SGD (batch size 1010). Results illustrate the impact of varying the number of clients ($M$), the number of local samples per client ($N_{i}$), and the heterogeneity parameter ($\epsilon$), with each client performing $K_{i}=2$ local updates at learning rate of $10^{-3}$. Sub-figures: (a) $N_{i}=10$, $\epsilon=0.01$; (b) $M=10$, $\epsilon=0.01$; (c) $M=10$, $N_{i}=10$.
Figure 5: Estimation error on synthetic data as a function of global iterations using gradient descent, evaluated across different client configurations of the number of clients ($M$), local dataset size per client ($N_{i}$), and heterogeneity parameter ($\epsilon$). In all cases, each client performs $K_{i}=5$ local update steps with a fixed learning rate of $10^{-4}$. The following configurations are considered: (a) $N_{i}=10$, $\epsilon=0.1$ with varying $M$; (b) $M=25$, $\epsilon=0.1$ with varying $N_{i}$; (c) $M=25$, $N_{i}=25$ with varying $\epsilon$.
Figure 3 demonstrates the experimental results for a real-world, nonlinear dynamical system a pendulum. Figure 3 demonstrates the substantial impact of client participation on estimation error. With a fixed number of trajectories per client ($N_{i}=10$) and low system heterogeneity ($\epsilon=0.01$), increasing the number of participating clients from 1 to 81 significantly reduces the estimation error. Figure 3 reveals the critical role of the data volume in convergence performance. When maintaining a fixed number of clients ($M_{i}=10$) and low system heterogeneity ($\epsilon=0.01$), increasing the number of trajectories per client accelerates error reduction. This scaling demonstrates how federated learning enables each client to benefit indirectly from the expanded data pool through collaborative learning. Figure 3 illustrates the adverse impact of system heterogeneity on convergence behavior. Under a fixed number of clients ($M_{i}=10$) and trajectories ($N_{i}=10$), increasing the heterogeneity parameter leads to higher estimation error. As $\epsilon$ grows, it introduces greater variability into the model, which in turn makes precise estimation more difficult and increases the overall error.
Figure 4 demonstrates the system behavior under stochastic gradient descent implemented with a batch size of 10 and two local updates per client. This configuration modifies both the granularity of the local optimization and the number of iterations per aggregation round. The introduction of minibatch stochastic gradient descent (SGD), rather than full-batch updates, further injects stochasticity into the optimization trajectory and may accelerate convergence and improve generalization on unseen trajectories. Using two local training updates enhances the local update effect; this can accelerate the reduction in estimation error under heterogeneous conditions, but may also exacerbate client drift in more heterogeneous settings. We systematically analyze gradient descent and mini-batch SGD for federated nonlinear system identification, offering practical insights for choosing algorithms under communication and computational constraints. Results for the quadrotor system have been provided in the Appendix.
Figure 5 presents empirical convergence results for federated nonlinear system identification conducted on synthetic datasets, where the state-input pairs are generated from a synthetic nonlinear dynamical system. Specifically, Figure 5 illustrates that increasing the number of clients enhances estimation accuracy, given a fixed number of trajectories per client $(N_{i}=25$) and a constant heterogeneity parameter $(\epsilon=0.1)$. Similarly, Figure 5 demonstrates that augmenting the number of trajectories per client improves estimation quality when both the number of clients $(M)$ and the heterogeneity parameter $(\epsilon=0.01)$ remain fixed. Conversely, Figure 5 indicates a degradation in estimation performance as the heterogeneity parameter increases, with the number of clients $(M)$ and trajectories per client $(N_{i})$ held constant. These observations collectively underscore the trade-offs between data distribution, quantity, and heterogeneity in federated nonlinear system identification.
Figure 6 shows graphs depicting how the normalized estimation error ($e$) varies with the square root of the number of clients ($\sqrt{M}$) for different values of epsilon ($\epsilon$) and fixed number of trajectories per client $(N_{i})$ for the pendulum dynamics and synthetic data. Figure 6 adopts a configuration analogous to that in Figure 3, whereas Figure 6 employs the setup outlined in Figure 5. This empirical validation highlights how the non-asymptotic convergence rate in federated learning can be systematically improved and scaled by a factor of $\sqrt{M}$ in the low heterogeneity regime.
Figure 6 illustrates that increasing the number of local updates per communication round can reduce communication overhead by lowering the frequency of global aggregations. However, beyond a certain point, the performance gains diminish as excessive local training in each global iteration introduces significant global model divergence across clients. The experimental setup for Figure 6 is similar to that of Figure 4, except for a variation in the number of local updates.
Figure 6: Comparison of estimation error versus $\sqrt{M}$ on (a) a real-world pendulum system and (b) synthetic data. The empirical results validate that, in low heterogeneity settings, the non-asymptotic convergence rate can be enhanced by increasing number of clients. (c) Impact of local updates $(K_{i})$ on estimation error.
6 Conclusion
We introduced a federated learning framework for identifying linearly parameterized nonlinear dynamical systems, with a particular focus on piecewise affine (PWA) models. Our theoretical analysis shows that the convergence error decreases as $1/\sqrt{M}$, where $M$ is the number of clients, thus having significant improvement in convergence as more clients collaborate. We corroborate our theory on the pendulum and quadrotor nonlinear physical systems, experimentally demonstrating the improvement in convergence error at any client decreases favorably as more clients collaborate for federated system identification.
Our work opens many interesting future directions. Theoretical analysis of how different optimization hyperparameters (e.g., number of local epochs in each global iteration) impact convergence is of significant interest. Adaptive batch size strategies could also balance the trade-off between gradient noise and convergence speed. A key fundamental open problem is where $\phi$ is not known a priori. One could explore scenarios where $\phi$ itself is learned collaboratively alongside the model, forming an end-to-end variant of our proposed framework. Finally, while this work focused on PWA models, a natural extension is to consider linearly parameterized nonlinear systems via Koopman theory (Brunton et al., 2022), which is based on lifting the states into an infinite-dimensional feature space, where the dynamics evolve linearly. While exact infinite-dimensional embeddings are impractical, one can explore learning a finite set of invariant eigenfunctions to approximate this transformation.
Appendix A Appendix
In this section, we provide definitions for the $\mathcal{K}$ and $\mathcal{KL}$ functions, proofs for Lemma 1, Proposition 1, Proposition 2, and Theorem 1. In the end, we provide the results for quadrotor experiments.
Definition 3 (Classes $\mathcal{K}$, and $\mathcal{KL}$).
The following function class definitions are cited from Kellett (2014).
  1. 1.
    A function $\alpha:[0,a)\to[0,\infty)$ (with $a\in(0,\infty]$) is of class $\mathcal{K}$ if it is continuous, strictly increasing, and $\alpha(0)=0$.
  2. 2.
    A function $\beta:[0,a)\times[0,\infty)\to[0,\infty)$ is of class $\mathcal{KL}$ if, for each fixed $t\geq 0$, the map $r\mapsto\beta(r,t)$ belongs to $\mathcal{K}$, and for each fixed $r\in[0,a)$, the map $t\mapsto\beta(r,t)$ is continuous, nonincreasing, and $\lim_{t\to\infty}\beta(r,t)=0$.
A.1 Proof of Lemma 1
Proof.
See Theorem 1 in Musavi et al. (2024). This lemma simply changes notation to fit the federated setting. ∎
A.2 Proof of Proposition 1
Proof.
Fix a client $i$ and a unit vector $v\in\mathbb{S}^{n_{\varphi}-1}$, and let $Y_{t}:=v^{\top}\varphi(x^{(i)}_{t},u^{(i)}_{t})$ for $t=0,\ldots,T_{i}-1$. By Lemma 1 the sequence $\{Y_{t}\}$ satisfies the $(1,s_{\varphi},p_{\varphi})$ block–martingale small-ball condition, so Proposition 2.5 of Simchowitz et al. (2018), with $k=1$ and $\alpha=1$, gives
$\Pr\!\bigl{[}\textstyle\sum_{t}Y_{t}^{2}\leq s_{\varphi}^{2}\!\bigl{(}N_{i}p_{\varphi}-2\log(1/\delta)\bigr{)}\bigr{]}\,\leq\,\delta.$
Take a $1/4$-net $\mathcal{N}$ of the unit sphere and apply the same bound to each $v\in\mathcal{N}$ with failure level $\delta/|\mathcal{N}|:=\delta$; a union bound gives that, with probability $\geq 1-\delta$, the inequality holds for every $v\in\mathcal{N}$. Because any unit $u$ is within $1/4$ of some $v\in\mathcal{N}$
$\lambda_{\min}\bigl{(}\Phi^{(i)}\Phi^{(i)\!\top}\bigr{)}\;\geq\;\frac{1}{2}\,s_{\varphi}^{2}\bigl{(}N_{i}Tp_{\varphi}-2\log(9^{n_{\varphi}}/\delta)\bigr{)}.$
If $N_{i}T\geq\tfrac{4}{p_{\varphi}}\bigl{[}n_{\varphi}\log 9+\log(1/\delta)\bigr{]}$, the component inside the bracket is at least $\tfrac{1}{2}N_{i}Tp_{\varphi}$, and replacing $p_{\varphi}$ by $1$ loses only another factor $2$, giving $\lambda_{\min}(\Phi^{(i)}\Phi^{(i)\!\top})\geq\tfrac{1}{2}s_{\varphi}^{2}N_{i}T$ with probability at least $1-\delta$. Finally, substituting $\delta/M$ for $\delta$ and union-bounding over the $M$ clients yields the solution. ∎
A.3 Proof of Proposition 2
Proof.
Conditioned on $\Phi^{(i)}$, the columns $w^{(i)}_{t}$ are independent, mean-zero, $\sigma_{w}$-sub-Gaussian vectors. For any $t=0,\dots,T-1$, the matrix-form Bernstein inequality gives
$\Pr\!\Bigl{(}\bigl{\|}W^{(i)}(\Phi^{(i)})^{T}\bigr{\|}_{2}>t\Bigr{)}\;\leq\;2\exp\!\Bigl{(}-\frac{t^{2}}{2\sigma_{w}^{2}\,\|\Phi^{(i)}\|_{F}^{2}}\Bigr{)}.$ (13)
Since $\|\phi(x,u)\|_{2}^{2}\leq b_{\phi}$, we have $\|\Phi^{(i)}\|_{F}^{2}\leq N_{i}T\,b_{\phi}\leq N_{i}T(n_{x}+n_{\phi})$. Choosing
$t\;=\;4\sigma_{w}\sqrt{\,N_{i}T\bigl{(}n_{x}+n_{\phi}+\log(2M/\delta)\bigr{)}}$
makes the right side of (13) less than or equal to $\delta/(2M)$. A union bound over $i\in[M]$ and some algebra completes the proof. ∎
A.4 Proof of Theorem 1
Proof.
Consider the relation for client $i$
$X^{(i)}_{+}\;=\;\theta^{(i)*}\,\Phi^{(i)}+W^{(i)}.$
Subtracting $\bar{\theta}_{\mathrm{LSE}}\Phi^{(i)}$ and regrouping gives
$X^{(i)}_{+}-\bar{\theta}_{\mathrm{LSE}}\Phi^{(i)}=(\theta^{(i)*}-\bar{\theta}_{\mathrm{LSE}})\,\Phi^{(i)}+W^{(i)}.$
Right-multiplying by $(\Phi^{(i)})^{\top}$ and summing over $i$,
$\sum_{i=1}^{M}(X^{(i)}_{+}-\bar{\theta}_{\mathrm{LSE}}\Phi^{(i)})(\Phi^{(i)})^{\top}=\sum_{i=1}^{M}(\theta^{(i)*}-\bar{\theta}_{\mathrm{LSE}})\,\Phi^{(i)}(\Phi^{(i)})^{\top}+\sum_{i=1}^{M}W^{(i)}(\Phi^{(i)})^{\top}.$
Defining [^2] Because $\bar{\theta}_{\mathrm{LSE}}$ minimizes the pooled least-squares objective, the left side sum vanishes. $G:=\sum_{i}\Phi^{(i)}(\Phi^{(i)})^{\top}$ and $\bar{\theta}^{*}:=\tfrac{1}{M}\sum_{i}\theta^{(i)*},$ we obtain
$G\,(\bar{\theta}_{\mathrm{LSE}}-\bar{\theta}^{*})^{\top}=-\sum_{i}(\theta^{(i)*}-\bar{\theta}^{*})\Phi^{(i)}(\Phi^{(i)})^{\top}-\sum_{i}W^{(i)}(\Phi^{(i)})^{\top}.$
We can bound each term as follows:
$G\succeq\tfrac{1}{2}\,s_{\phi}^{2}N_{\mathrm{tot}}\,I$
by Proposition 1,
$\sum_{i}^{M}W^{(i)}(\Phi^{(i)})^{T}\leq 4\sigma_{w}\sqrt{N_{i}T(n_{x}+n_{\phi}+log(2M/\delta))}$
by Proposition 2, and
[…]
using Assumption 5  [^3] Notice that Assumption 5 implies that $\theta^{(i)*}-\bar{\theta}^{(i)*}||\leq\epsilon$, $\forall i\in[M]$. and that $\Phi^{(i)}(\Phi^{(i)})^{\top}||_{2}\leq N_{i}b_{\phi}$.
By multiplying both sides by $G^{-1}$, applying the bounds, taking the norm of each side gives with probability greater than or equal to $1-3\delta$,
$\bar{\theta}_{LSE}-\bar{\theta}^{*}||_{2}\leq\frac{8\sigma_{w}}{s_{\phi}^{2}}\sqrt{\frac{n_{x}+n_{\phi}+log(2M/\delta)}{\sum_{i=1}^{M}N_{i}T}}+\frac{b_{\phi}}{s_{\phi}^{2}}\epsilon.$
Since Assumption 5 implies $\|\theta^{(i)*}-\bar{\theta}^{*}\|_{2}\leq\varepsilon/2$ for $i\in[M]$,
$\|\bar{\theta}_{\mathrm{LSE}}-\theta^{(i)*}\|_{2}\leq\|\bar{\theta}_{\mathrm{LSE}}-\bar{\theta}^{*}\|_{2}+\varepsilon/2.$
This yields the expected bound. ∎
Figure 7: Estimation error versus the number of global iterations for the real-world nonlinear dynamical system of a quadrotor using gradient descent (GD). Results illustrate the impact of varying the number of clients ($M$), the number of local samples per client ($N_{i}$),with each client performing $K_{i}=5$ local updates at a learning rate of $10^{-1}$. Sub-figures: (a) $N_{i}=10$, $\epsilon=2$; (b) $M=5$, $\epsilon=5$.
A.5 Results for quadrotor
Another real-world system considered is the quadrotor, adapted from the model proposed in Musavi et al. (2024); Alaimo et al. (2013), with necessary modifications as described in the following. Let $p\in\mathbb{R}^{3}$ and $v\in\mathbb{R}^{3}$ denote the position and velocity of the quadrotor’s center of mass in the inertial frame, respectively. Let $\omega\in\mathbb{R}^{3}$ represent the angular velocity in the body-fixed frame, and $q\in\mathbb{R}^{4}$ represent the orientation of the quadrotor using a unit quaternion. The corresponding equations of motion for the quadrotor are given by:
$\dfrac{d}{dt}\begin{pmatrix}p\\ v\\ q\\ \omega\end{pmatrix}=\begin{pmatrix}v\\ -ge_{z}+\frac{1}{m}Qf_{u}\\ \frac{1}{2}\Omega q\\ I^{-1}(\tau_{u}-\omega\times I\omega)\end{pmatrix}+w,$
where $g$ is the gravitational constant, $m$ is the total mass of the quadrotor, $I=\emph{diag}(I_{xx},I_{yy},I_{zz})$ denotes the inertia matrix in the body-fixed frame, $f_{u}\in\mathbb{R}$ is the total thrust, $\tau_{u}\in\mathbb{R}^{3}$ is the total moment expressed in the body-fixed frame, and $e_{z}=(0,0,1)^{\top}$ is the unit vector along the inertial $z$-axis. There are seven unknown parameters in total, with heterogeneity introduced in a subset of these parameters.
$\theta_{1}$ $=\frac{1}{m}+\gamma_{1}^{(i)},\quad\theta_{2}=\frac{1}{I_{xx}},\quad\theta_{3}=\frac{I_{yy}-I_{zz}}{I_{xx}},$
$\theta_{4}$ $=\frac{1}{I_{yy}},\quad\theta_{5}=\frac{I_{zz}-I_{xx}}{I_{yy}},\quad\theta_{6}=\frac{1}{I_{zz}},$
$\theta_{7}$ $=\frac{I_{xx}-I_{zz}}{I_{zz}}.$
Here $\gamma_{1}^{(i)}\sim\mathrm{U}(0,\epsilon)$.
Figure 7 demonstrates the experimental results for quadrotor. Figure 7(a) demonstrates the impact of client participation on estimation error. With a fixed trajectory count per client ($N_{i}=10$) and system heterogeneity ($\epsilon=2$), increasing the number of participating clients reduces the error. Figure 7(b) underscores the role of data volume in convergence. Keeping the number of clients fixed ($M_{i}=5$) and heterogeneity ($\epsilon=5$) as fixed, increasing trajectories per client significantly accelerates error reduction. The length of each trajectory is set to 10 timesteps. These results follow similar trends as seen in the experiments for pendulum and synthetic datasets presented in the main paper.

### v3

@@
-Federated Nonlinear System Identification
+arXiv:2508.15025v3[cs.LG] 27 Feb 2026
 Abstract
@@
 footnotetext: * denotes equal contribution
-1 Introduction
+I Introduction
 Dynamical system models describe how a system evolves over time based on its current state, control inputs, and external disturbances.
-These models are central to fields including control theory, physics, and robotics (Willems, 1989).
-A common class of such systems is linear time-invariant (LTI) systems, where the next state and output are linear functions of the current state and input, governed by parameters that remain constant over time (Sibai & Mitra, 2018).
-In contrast, nonlinear systems allow these dependencies to be governed by nonlinear functions, offering more accurate representations for many real-world systems with complex or nonstationary behavior (Guckenheimer & Holmes, 2013; Jiang & Lai, 2019).
+These models are important to fields such as control theory, physics, and robotics [1].
+A common category of systems is linear time-invariant (LTI) systems, where both the next state and output are linear functions of the current state and input.
+These systems are governed by parameters that remain constant over time [2].
+In contrast, nonlinear functions determine these relationships in nonlinear systems [3, 4].
 Figure 1: Federated learning framework for nonlinear dynamical system identification involving $M$ clients which are similar but non-identical in nature and a central server.
@@
 The server then aggregates the models to obtain an updated global model for the next round.
-System identification is the process of learning dynamical models from input and output data.
-In this way, one can learn the dynamics of a system in a data-driven way, without explicit physical derivations.
-It is often used to learn dynamical systems in the real world, where exact dynamics are unknown, with applications including robotics, fluid dynamics, and physics (Sarkar et al., 2021; Venkatesh & Dahleh, 2002).
-Traditional approaches for system identification typically assume access to centralized datasets.
-However, in modern deployments—such as multi-robot systems, sensor networks, and connected vehicles—data is naturally distributed across many devices.
-Transmitting all data to a central location is often impractical due to privacy, bandwidth, or energy limitations.
+System identification is a data-driven approach to learn dynamical models from input and output data.
+It is often used to learn dynamical systems for which exact dynamics are unknown, with applications including robotics, fluid dynamics, physics and healthcare [5, 6, 7].
+Traditional approaches for system identification assume centralized data.
+However, in modern applications, data is often distributed across many devices, and transmitting all data to a central server is often impractical due to privacy, bandwidth, or energy limitations.
 To address these challenges, we consider federated system identification, which is emerging as a privacy-preserving alternative.
 In this paradigm, multiple clients collaboratively learn a model without sharing raw data, communicating only model updates.
-Prior work has explored federated identification for LTI systems (Wang et al., 2023), demonstrating convergence benefits that scale with the number of participating clients.
-However, this line of work is restricted to linear models, which fundamentally limits its applicability.
+Prior work has explored federated identification for LTI systems [8].
 In this paper, we study the problem of federated nonlinear system identification.
-We consider a setting where $M$ distributed clients each observe trajectory data from potentially different nonlinear dynamical systems that nonetheless belong to the same underlying family (see Figure 1).
-Dynamical systems at clients exhibit heterogeneity, modeled via a heterogeneity parameter $\epsilon$.
-This setup captures realistic scenarios such as learning dynamics across a fleet of heterogeneous robots or drones—each operating under different conditions due to hardware differences, controller tunings, or external disturbances.
-The goal is to identify a shared nonlinear model that generalizes across these clients while respecting local variability and preserving data privacy.
-1.1 Key Contributions
-Figure 2: Impact of client count and heterogeneity on federated nonlinear system identification.
-(a) Estimation error versus number of clients $M$ at fixed trajectories per client $N_{i}=10$ and heterogeneity bound $\epsilon=0.01$; (b) Estimation error versus $M$ for the same $N_{i}$ with $\epsilon=5$, illustrating increased model-mismatch tolerance; (c) Estimation error as a function of $\epsilon$ for fixed $M$ and $N_{i}$, showing sensitivity of global model accuracy to client-level parameter variations.
+We consider a setting where $M$ distributed clients each observe trajectory data from different nonlinear dynamical systems that belong to the same underlying family (see Figure 1).
+The dynamical systems we consider exhibit heterogeneity, modeled by the heterogeneity parameter $\epsilon$.
+Federated learning should be used for system identification when multiple similar systems share common structure and data cannot be centralized, enabling privacy-preserving collaborative learning.
+Results show that aggregating across more participating clients improves convergence and reduces individual client noise.
+Our main contributions are as follows:
   • Nonlinear framework.
-We consider linearly parameterized nonlinear dynamical systems, with a focus on the widely used piecewise affine (PWA) model for representing nonlinear dynamics (Mania et al., 2022; Musavi et al., 2024).
+We consider linearly-parameterized nonlinear dynamical systems, with a focus on the widely used piecewise affine (PWA) model for representing nonlinear dynamics [9, 10].
 In PWA models, state transitions are expressed as linear functions of nonlinear feature embeddings of state-input pairs.
 While prior work has explored federated system identification of linear dynamical systems, to the best of our knowledge, this work develops the first federated learning framework for system identification of PWA-based nonlinear dynamical systems.
-  • Theoretical analysis.
-We theoretically analyze the convergence of the federated nonlinear system identification problem.
-Our results show that clients benefit significantly by collaborating through federated learning.
-Particularly, we show that the convergence error decreases as $\mathcal{O}(1/\sqrt{M})$, where $M$ is the number of clients, thus having significant improvement in convergence as more clients collaborate.
-Additionally, our results theoretically characterize how heterogeneity across clients impacts the convergence error.
+[…]
   • Experimental analysis.
-To corroborate our theoretical results, we conduct extensive experiments analyzing the convergence behavior of our approach as a function of the total number of participating clients, the number of local samples at each client, as well the heterogeneity of dynamical systems across clients.
-We preview some results in Figure 2, which are based on real-world pendulum dynamics.
-Complete details are provided in Section 4.
-  • Algorithm validation.
-Our experiments demonstrate that our federated approach achieves improved convergence rate compared to a single-client system, while offering the additional advantage of client data privacy.
-2 Related Works
-2.1 System Identification
-Ljung (1998) provides an overview of mathematical methods for building models of dynamical systems using input and output signal measurements, focusing on how parameter estimation techniques can be applied to characterize system behavior in both time and frequency domains.
+We conduct experiments analyzing the convergence behavior of our approach as a function of the total number of participating clients, the number of local samples at each client, as well the heterogeneity of dynamical systems across clients.
+Complete details are provided in Section IV.
+  • Algorithm validation. Our experiments demonstrate that our federated approach achieves an improved convergence rate compared to a single-client system.
+II Related Works
+II-A System Identification
+The paper [11] provides an overview of mathematical methods for building models of dynamical systems using input and output signal measurements, focusing on how parameter estimation techniques can be applied to characterize system behavior in time and frequency domains.
 This approach is instrumental in transforming observed data into predictive models, supporting applications in signal processing and control.
-Jedra & Proutiere (2020) present a robust finite-time error bound for identifying stable linear systems using the ordinary least-squares (OLS) estimator.
-Simchowitz et al.
-(2018) also analyze OLS identification from a single trajectory for stable and marginally stable linear time-invariant systems, using the block-martingale small-ball method to address mixing dependencies and establish nearly optimal estimator rates for a wide class of linear time series.
-While deriving tight, regime-specific finite-time error rates for OLS identification of general linear systems, Sarkar & Rakhlin (2019) establish that OLS is optimally consistent only under specific regularity conditions, unifying previous results across all spectral regimes.
-However, this approach struggles with nonlinear or complex signal structures, often requiring significant computational resources and large datasets to achieve reliable identification.
-2.2 Federated Learning
-Federated learning is a machine learning paradigm which allows clients to collaboratively train a model without sharing raw data (McMahan et al., 2017).
-In principle, each client possesses a local dataset and sends weight updates to the server.
-The server then aggregates these weights together to update the global model.
-This approach is very useful to train models on sensitive or private data.
-Truex et al.
-(2019) propose a hybrid federated learning approach that combines differential privacy and secure multiparty computation to enhance privacy protection while maintaining high model accuracy in collaborative machine learning settings.
-The paper presents communication-efficient federated learning strategies through structured low-dimensional updates and compressed sketched updates, achieving a substantial reduction in communication overhead while preserving model accuracy (Konečnỳ et al., 2016).
-Li et al.
-(2020) introduce the FedProx framework, which adds a proximal term to FedAvg and provides convergence guarantees to achieve more robust and stable training in the presence of statistical and system heterogeneity across clients.
-Federated learning architectures can use a peer-to-peer approach (Yang et al., 2019).
-However, our discussion centers on the traditional server-client model, in which a central server periodically aggregates updates from multiple clients.
-While most federated learning works are concerned with learning deep neural networks, our work focuses on understanding the theory and implementation of federated system identification of nonlinear dynamical systems.
-2.3 Federated System Identification
-The goal of the federated system identification problem is to determine a shared estimate of the system matrices, denoted as $\bar{\theta}$, that yields a small estimation error relative to each client $i$’s true system parameters $\theta^{(i)*}$.
+System identification can be extended to linearly-parameterized nonlinear systems [10].
+The authors in [12] present a robust finite-time error bound for identifying stable linear systems using the ordinary least-squares (OLS) estimator.
+The study in [13] also analyzes OLS identification from a single trajectory for stable and marginally stable linear time-invariant systems, using the block-martingale small-ball method to address mixing dependencies and establishes nearly optimal estimator rates for a wide class of linear time series.
+The work further provides tight, regime-specific finite-time error bounds for OLS identification of general linear dynamical systems.
+The framework in [14] establishes that OLS is optimally consistent only under specific regularity conditions, unifying previous results across all spectral regimes.
+This approach struggles with nonlinear or complex signal structures, often requiring significant computational resources and large datasets to achieve reliable identification.
+II-B Federated Learning
+Federated learning is a machine learning paradigm which allows clients to collaboratively train a model while preserving data privacy [15].
+Each client possesses a local dataset and sends weight updates to the server.
+The server then aggregates these weights and updates the global model.
+The method presented by [16] is a hybrid federated learning approach that combines differential privacy and secure multi-party computation to enhance privacy protection while maintaining high model accuracy.
+Communication-efficient federated learning strategies through structured low-dimensional updates and compressed sketched updates may achieve a substantial reduction in communication overhead while preserving model accuracy [17].
+The work of[18] introduces the FedProx framework, which adds a proximal term to FedAvg and provides convergence guarantees to achieve more robust and stable training in the presence of statistical and system heterogeneity across clients.
+Federated learning architectures can use a peer-to-peer approach [19], however, our discussion focuses on the traditional server-client model, in which a central server periodically aggregates updates from multiple clients.
+Although most federated learning works are concerned with learning deep neural networks, our work focuses on understanding the theory and implementation of federated system identification of nonlinear dynamical systems.
+II-C Federated System Identification
+The goal of federated system identification is to learn a shared estimate of the system parameters that yields a small estimation error relative to the true parameters of each client $i$’s true system parameters $\theta^{(i)*}$.
 Clients $i$ do not share their private data with the server, and each client independently estimates its unknown system dynamics.
-Wang et al.
-(2023) formulated an FL framework for system identification of linear system models across heterogeneous clients with improved sample efficiency.
-They show that more clients improve the model convergence rate and propose the FedSysID algorithm.
-Their work was restricted to synthetic linear systems, which limits applicability to real-world nonlinear dynamical systems.
-This limitation motivates our proposal of federated nonlinear system identification, as detailed in the following sections.
-3 Federated Nonlinear System Identification
+The paper [8] formulated an FL framework for system identification of linear system models across heterogeneous clients with improved sample efficiency, proposing the FedSysID algorithm and showing that more clients improve the model convergence rate.
+However, [8] was restricted to synthetic linear systems.
+The authors of [20] introduce Koopman learning as a method for dealing with nonlinear systems that have partially observable data.
+The setting is where multiple clients observe an identical nonlinear system, but the system lacks any active control ($u_{t})$, which is contrary to what we will consider in (1).
+These limitations motivate our work on federated nonlinear system identification.
+III Federated Nonlinear System Identification
 In this section, we formally define the federated nonlinear system identification framework.
@@
 Finally, we present the FNSysId algorithm that describes the client-server interactions for nonlinear system identification.
-3.1 Preliminaries
+III-A Preliminaries
 The formal objective of the federated system identification problem is to estimate a model that accurately captures the collective behavior of the participating systems.
-Specifically, suppose that for each client $i\in[M]$, we have the following linearly parameterized nonlinear dynamics:
+Formally, first suppose that for each client $i\in[M]$, we have the following linearly-parameterized nonlinear dynamics:
 $x_{t+1}^{(i)}=\theta^{(i)*}\phi(x_{t}^{(i)},u_{t}^{(i)})+w_{t}^{(i)},$ (1)
 where $x_{t}^{(i)}$, $u_{t}^{(i)}$, and $w_{t}^{(i)}$ denote the state, control, and disturbance respectively.
-Furthermore, $\phi:\mathbb{R}^{n_{x}}\times\mathbb{R}^{n_{n}}\to\mathbb{R}^{n_{\phi}}$ denotes a vector of known nonlinear mappings (Mania et al., 2022; Musavi et al., 2024).
-In the following, we recall two key definitions from Musavi et al. (2024) that will be used in our assumptions.
-Definition 1.
-(Semi-continuous distribution). A probability distribution $\mathbb{P}$ is semi-continuous if there does not exist a set $E$ with Lebesgue measure zero such that $\mathbb{P}(E)=1$.
-Definition 2.
+Furthermore, $\phi:\mathbb{R}^{n_{x}}\times\mathbb{R}^{n_{u}}\to\mathbb{R}^{n_{\phi}}$ denotes a vector of known nonlinear mappings [9, 10].
+In the following, we recall two key definitions from [10] that will be used in our assumptions. Refer to [10] for further clarification on these definitions.
+Definition 1
+(Semi-continuous distribution). A Euclidean probability distribution $\mathbb{P}$ is semi-continuous if there does not exist a set $E$ with Lebesgue measure zero such that $\mathbb{P}(E)=1$.
+Definition 2
 […]
-We next provide four assumptions for the local dynamical systems at clients (Musavi et al., 2024) (we omit client indexing for brevity):
+We next provide four assumptions for the local dynamical systems at clients [10] (we omit client indexing for brevity).
   • Assumption 1: All components of the feature vector $\phi(\cdot)$ are real-analytic functions in $\mathbb{R}^{n_{x}+n_{u}}$, i.e., all components in $\phi(\cdot)$ are infinitely differentiable.
-  • Assumption 2: Noise $w_{t}$ is i.i.d., and follows a semi-continuous, mean-zero distribution with positive semi-definite covariance matrix $\sum_{w}\succeq\sigma_{w}^{2}I_{n_{x}}$.
-  • Assumption 3: Input $u_{t}$ is i.i.d following a semi-continuous, mean-zero distribution with positive semi-definite covariance matrix $\sum_{u}\succeq\sigma_{u}^{2}I_{n_{x}}$.
+  • Assumption 2: Noise $w_{t}$ is i.i.d., and follows a semi-continuous, zero-mean distribution with positive semi-definite covariance matrix $\sum_{w}\succeq\sigma_{w}^{2}I_{n_{x}}$.
+  • Assumption 3: Input $u_{t}$ is i.i.d following a semi-continuous, zero-mean distribution with positive semi-definite covariance matrix $\sum_{u}\succeq\sigma_{u}^{2}I_{n_{x}}$.
 […]
-Additionally, inspired by Wang et al. (2023), we consider the following assumption of bounded system heterogeneity across clients.
+Additionally, inspired by [8], we consider the following assumption of bounded system heterogeneity across clients.
 […]
@@
 Functionally, this optimization estimates each client’s local dynamics via least-squares regression, then averages the results to obtain a global model.
-3.2 Framework Convergence Rate
+III-B Convergence Rate
 A least-squares estimate is meaningful only if the design matrix built from each client’s regressor vectors is well conditioned.
-In linear system identification this role is played by a persistent-excitation assumption.
-For linearly parameterized nonlinear dynamics, we recover a similar property through the block-martingale small-ball (BMSB) condition.
+In linear system identification this role is played by a persistent‐excitation assumption.
+For linearly-parameterized nonlinear dynamics, we recover a similar property through the block‑martingale small‑ball (BMSB) condition (Definition 4).
 Lemma 1 is used to show that the analytic feature map and new inputs guarantee the required excitation.
-Lemma 1 (BMSB for open-loop systems).
-Let each client $i\in[M]$ run open-loop inputs $u^{(i)}_{t}=\eta^{(i)}_{t}$. Let client $i$ collect a trajectory of length $T$. For every $t=0,1,\dots,T-1$, define the filtration
+Lemma 1 (BMSB for open‑loop systems)
+Let each client $i\in[M]$ run open‑loop inputs $u^{(i)}_{t}=\eta^{(i)}_{t}$. Let client $i$ collect a trajectory of length $T$. For every $t=0,1,\dots,T-1$, define the filtration
 $\mathcal{F}^{(i)}_{t}:=\sigma\bigl(w^{(i)}_{0},\dots,w^{(i)}_{t-1},\;x^{(i)}_{0},\dots,x^{(i)}_{t},\;\eta^{(i)}_{0},\dots,\eta^{(i)}_{t}\bigr)$.
-Under Assumptions 1–4 there exist constants $s_{\phi}>0$ and $p_{\phi}\in(0,1)$ (as defined by Musavi et al.
-(2024)) such that, for every client $i$ and every unit vector $v\in\mathbb{S}^{n_{\phi}-1}$,
+Under Assumptions 1–4 there exist constants $s_{\phi}>0$ and $p_{\phi}\in(0,1)$ (as defined by [10]) such that, for every client $i$ and every unit vector $v\in\mathbb{S}^{n_{\phi}-1}$,
 $\Pr\!\bigl(|v^{\!\top}\phi(x^{(i)}_{t},u^{(i)}_{t})|\,\geq s_{\phi}\;\big|\;\mathcal{F}^{(i)}_{t-1}\bigr)\;\geq\;p_{\phi}\quad\text{a.s.}$ (3)
 Hence the regressor process $\{\phi(x^{(i)}_{t},u^{(i)}_{t})\}_{t\geq 0}$ satisfies the $(1,s_{\phi}^{2}I_{n_{\phi}},p_{\phi})$-BMSB condition for every client $i$.
-Proof.
+Proof:
 Please refer to the Appendix for the proof. ∎
-Less formally, each new regressor $\phi(x^{(i)}_{t},u^{(i)}_{t})$ has (conditioned on the past) at least probability $p_{\phi}$ of having length $s_{\phi}$ in every direction, hence the excitation we desired.
+Each new regressor $\phi(x^{(i)}_{t},u^{(i)}_{t})$ has (conditioned on the past) at least probability $p_{\phi}$ of having length $s_{\phi}$ in every direction, hence the excitation we desired.
 To use least-squares, the empirical Gram matrices must be invertible. Proposition 1 provides that guarantee and will be used to determine the $1/N_{\mathrm{tot}}$ factor in our final error bound.
-Proposition 1.
+Proposition 1
 Fix $\delta\in(0,1)$.
@@
 $G:=\sum_{i=1}^{M}\Phi^{(i)}\Phi^{(i)\!\top}\succeq\tfrac{1}{2}\,s_{\phi}^{2}N_{\mathrm{tot}}\,I_{n_{\phi}}.$ (4)
-Proof.
+Proof:
 Please refer to the Appendix for the proof. ∎
 We now need to bound the stochastic noise regressor cross-term. This will be used to guarantee that the estimator remains stable and accurate despite disturbances.
-Proposition 2.
+Proposition 2
 Fix $0<\delta<1$. For client $i\in[M]$ define the noise matrix
@@
 $\|P\|_{2}\;\leq\;4\sigma_{w}\sqrt{\,N_{i}T(n_{x}+n_{\phi}+\log(2M/\delta))}.$ (6)
-Proof.
+Proof:
 Please refer to the Appendix for the proof. ∎
 These auxiliary propositions can now be combined to bound the final convergence rate.
-Theorem 1 (Finite-sample error).
+Theorem 1 (Finite‑sample error)
 Let $s_{\phi}$, $p_{\phi}$ be as in Lemma 1 and set
@@
 for all $i\in[M].$
-Proof.
+Proof:
 Please refer to the Appendix for the proof. ∎
-Remark 1.
+Remark 1
 The first term in the convergence error decreases as $1/\sqrt{T\sum_{i=1}^{M}N_{i}}$, where $\sum_{i=1}^{M}N_{i}$ is the total number of trajectories across all clients.
-Furthermore, since the number of trajectories $N_{i}$ at client $i$ is a constant for a given trajectory length $T$, the error decreases as $\mathcal{O}(1/\sqrt{M})$.
-Remark 2.
+Furthermore, since the number of trajectories $N_{i}$ at client $i$ is a constant for a given trajectory length $T$, the error decreases as $\tilde{\mathcal{O}}(1/\sqrt{M})$.
+Remark 2
 For smaller $\epsilon$, the underlying system dynamics at the clients are more similar, enabling FL clients to effectively use data from all participants.
-This enhances the performance of each client, particularly achieving a convergence speedup of $\sqrt{M}$.
-3.3 FNSysId
-The core workflow of FNSysId implementation is formalized in Algorithm 1, which shows how clients perform local computations and synchronize model updates with the central server during each communication round.
-Algorithm 1 FNSysId
-Initialize the server with $\bar{\theta}_{0}$, and learning rate $\alpha$ ;
-Initialize each client $i\in[M]$ with $\theta_{0,0}^{{(i)}}=\bar{\theta}_{0}$;
-For each global iteration $r=0,1,\ldots,R-1$ do
-[…]
-  For each client $i\in[M]$ in parallel do
-   $\theta_{r+1}^{{(i)}}=\texttt{ClientUpdate}(i,\bar{\theta}_{r},K_{i})$
-   send $\theta_{r+1}^{{(i)}}$ back to the server
-  end for
-[…]
-  update $\bar{\theta}_{r+1}=\frac{1}{M}\sum_{i=1}^{M}\theta_{r+1}^{{(i)}}$ and send $\bar{\theta}_{r+1}$ to each client
-end for
-Return $\bar{\theta}_{R}$
+This enhances the performance of each client, particularly achieving a convergence speedup of ${\approx}\sqrt{M}$.
+III-C FNSysId
+The FNSysId  [^2] The source code is available for download - https://github.com/OMKARTT/FNSysId implementation is formalized in Algorithm 1, which shows how clients perform computations and synchronize model updates with the central server during each communication round.
+This is the standard FL algorithm.
 Each client initializes a guess for its local model parameters, denoted as $\bar{\theta}_{0}$, and a learning rate $\alpha$.
@@
 In our framework, we assume all clients $\mathcal{C}=\{C_{1},C_{2},\ldots,C_{M}\}$ actively participate in every global iteration $r\in[R]$.
-Each client performs $K_{i}$ local updates independently using its own local data before any communication with the central server (line 6), where ClientUpdate iteratively executes:
-$\theta_{r,k}^{{(i)}}=\;$ $\theta_{r,k-1}^{{(i)}}+\alpha\left({y}^{{(i)}}-\theta_{r,k-1}^{{(i)}}{X}^{{(i)}}\right)({X}^{{(i)}})^{\top},\quad k=1,2,\ldots,K_{i}.$ (7)
+Each client performs $K_{i}$ local updates independently using its own local data before any communication with the central server (line 6), where NonlinearClientUpdate iteratively executes:
+$\theta_{r,k}^{{(i)}}=\;$ $\theta_{r,k-1}^{{(i)}}+\alpha\left({y}^{{(i)}}-\theta_{r,k-1}^{{(i)}}{X}^{{(i)}}\right)(X^{(i)})^{T},$ (7)
+where $X^{(i)}$ is the input data and $y^{(i)}$ is the true system. $k=1,2,\ldots,K_{i}.$ Note that this is a restatement of FedAvg.
 This approach reduces communication overhead by allowing multiple updates locally.
 After completing these updates, the clients send their updated local models $\bar{\theta}_{r+1}^{(i)}$ to the server.
-The server then averages all updates to obtain a new global model $\bar{\theta}_{r+1}$, which is subsequently broadcast back to all clients to synchronize (line 10).
+The server then averages all updates to obtain a new global model $\bar{\theta}_{r+1}$, which is subsequently broadcast back to all clients to synchronize (line 11).
 This efficiently balances local computation and communication, addressing potentially limited communication capabilities between clients.
-4 Experiments
-We conducted experiments on a range of systems, encompassing both synthetic benchmarks and real-world dynamical systems.
-In the following, we provide details regarding datasets employed, hyperparameter configurations, and the corresponding results.
-4.1 Data
-Conducting experiments on both synthetic and real-world system data is essential to comprehensively evaluate the proposed methodology.
-Synthetic datasets offer controlled environments enabling precise assessment of model accuracy and convergence behavior.
-In contrast, real-world datasets provide insights into the method’s practical applicability, robustness to noise, and generalization capabilities under complex, uncontrolled conditions.
-This dual evaluation ensures both theoretical soundness and empirical relevance.
-4.1.1 Synthetic Data
-To evaluate the performance and efficiency of FNSysId, we conduct numerical experiments involving $M$ clients, each associated with a discrete-time dynamical system characterized by $n=3$ states and $p=2$ inputs.
+Algorithm 1 FNSysId: Federated Nonlinear System Identification
+1:Initialize server with global parameter $\bar{\theta}_{0}$ and learning rate $\alpha$ ;
+2:Initialize each client $i\in[M]$ with local parameter $\theta^{(i)}_{0,0}=\bar{\theta}_{0}$ ;
+3:for global iteration $r=0,1,\ldots,R-1$ do
+[…]
+5:  for each client $i\in[M]$ in parallel do
+6:    $\theta^{(i)}_{r+1}=\texttt{NonlinearClientUpdate}\big(i,\bar{\theta}_{r},K_{i}\big)$ (7)
+7:    Send $\theta^{(i)}_{r+1}$ to server
+8:  end for
+[…]
+10:  $\bar{\theta}_{r+1}=\frac{1}{M}\sum_{i=1}^{M}\theta^{(i)}_{r+1}$
+11:  Broadcast $\bar{\theta}_{r+1}$ to all clients
+12:end for
+13:Return global estimate $\bar{\theta}_{R}$
+Corollary 1
+Using the same setup as Theorem 1 and for all communication rounds $R\geq 1$, the output $\bar{\theta}_{R}$ given by the FNSysId algorithm with the FedAvg update rule satisfies:
+$\max_{i\in[M]}\|\bar{\theta}_{R}-\theta^{(i)*}\|_{2}\leq\mathcal{O}\left(\frac{1}{KR}+\frac{1}{\sqrt{\sum_{i=1}^{M}N_{i}T}}+\epsilon\right).$
+Proof:
+Because the system studied in this paper is linearly-parameterized, the proof directly follows from [8, Corollary 4]. ∎
+Figure 2: Estimation error versus the number of global iterations for the real-world nonlinear dynamical system of a pendulum using gradient descent (GD).
+Results illustrate the impact of varying the number of clients ($M$), the number of local samples per client ($N_{i}$), and the heterogeneity parameter ($\epsilon$), with each client performing $K_{i}=1$ local updates at alearning rate of $10^{-2}$.
+Subfigures: (a) $N_{i}=10$, $\epsilon=0.01$; (b) $M=10$, $\epsilon=0.01$; (c) $M=20$, $N_{i}=10$.
+Figure 3: Estimation error on synthetic data as a function of global iterations using gradient descent, evaluated across different client configurations of the number of clients ($M$), local dataset size per client ($N_{i}$), and heterogeneity parameter ($\epsilon$).
+In all cases, each client performs $K_{i}=5$ local update steps with a fixed learning rate of $10^{-4}$.
+The following configurations are considered: (a) $N_{i}=10$, $\epsilon=0.1$ with varying $M$; (b) $M=25$, $\epsilon=0.1$ with varying $N_{i}$; (c) $M=25$, $N_{i}=25$ with varying $\epsilon$.
+IV Experiments
+IV-A Data
+In this section we describe the datasets used in our experiments.
+IV-A1 Synthetic Data
+To evaluate the performance and efficiency of FNSysId, we conduct experiments involving $M$ clients, each associated with a discrete-time dynamical system characterized by $n=3$ states and $p=2$ inputs.
 The initial state, input signal, and process noise are all sampled from zero-mean distributions with unit standard deviations.
-The trajectories were generated by augmenting the trajectory formulation presented by Wang et al.
-(2023) with a nonlinear term, as presented in the following equation.
+The trajectories were generated by augmenting the trajectory formulation presented by [8] with a nonlinear term, as presented in the following equation.
 $x_{t+1}^{(i)}=A^{(i)}\sin(x_{t}^{(i)})+B^{(i)}u_{t}^{(i)}+w_{t}^{(i)}.$ (8)
-Following the approach of Xin et al.
-(2022), we construct heterogeneous client dynamics by perturbing a nominal system $(A_{0},B_{0})$.
-Particularly, for client $i$, the perturbed system matrices are as follows:
+Following the approach of [21], we construct heterogeneous client dynamics by perturbing a nominal system $(A_{0},B_{0})$. Particularly, for client $i$, the perturbed system matrices are as follows:
 $A^{(i)}=A_{0}+\gamma_{1}^{(i)}V,$ (9)
@@
 We consider a trajectory length of $T=5$ and assume that each client uses an equal number of data points.
-4.1.2 Real-World System
-We next conduct experiments across two different physical systems whose characteristics and control requirements provide contrasting evaluation scenarios.
-The first is a pendulum system characterized by one state and one input, targeting the estimation of two unknown parameters.
-Further details on the trajectory generation methodology are provided in Section 4 of Musavi et al.
-(2024).
-In all experimental settings, the control input $u_{t}$ is defined as the superposition of a deterministic policy and stochastic noise, i.e., $u_{t}=\pi(x_{t})+\eta_{t},$ where $\pi(x_{t})$ denotes the control policy evaluated at state $x_{t}$ and $\eta_{t}$ represents zero-mean i.i.d.
-noise.
-Here $\eta_{t}$ is sampled from uniform distribution distribution to model bounded stochastic disturbances.
-For the pendulum, we have two nominal parameters: $A_{0}=\tfrac{1}{l}$ and $B_{0}=\tfrac{1}{ml^{2}}$.
-Here, $m$ denotes the mass of the pendulum and $l$ represents its length.
-Following the formulation of Musavi et al.
-(2024), we modified the trajectories by introducing the heterogeneity using the same procedure outlined in equation 9 and equation 10.
+IV-A2 Real-World System
+We conduct experiments across two different physical systems.
+The pendulum system is characterized by one state and one input, targeting the estimation of two unknown parameters [^3] Details on the trajectory generation method are in [10, Section 4].
+In all both experiments, the control input is defined as $u_{t}=\pi(x_{t})+\eta_{t},$ where $\pi(x_{t})$ denotes the control policy evaluated at state $x_{t}$ and $\eta_{t}$ represents zero-mean i.i.d.
+noise, sampled.
+Here $\eta_{t}$ is sampled from a uniform distribution.
+We have two nominal parameters: $A_{0}=\tfrac{1}{l}$ and $B_{0}=\tfrac{1}{ml^{2}}$ ($m$ denotes the mass of the pendulum and $l$ represents its length).
+Following the formulation of [10], we modified the trajectories by introducing the heterogeneity using the same procedure outlined in equations (9) and (10).
 $\ddot{\alpha}=-A^{(i)}g\sin(\alpha)+B^{(i)}u+w,$ (11)
-where $\alpha$ represents the angle of the rod relative to the vertical axis, $g$ is the gravity constant, $u$ is the torque input, and $w$ is the disturbance applied to this system.
-After discretization, the system dynamics can be rewritten in the structure of (1) with the feature vector consisting of expressions involving $\sin(\alpha)$ and $u$, all of which are analytic functions.
-The matrix of unknown parameters contains terms of the pendulum’s mass and the rod’s length.
-Each trajectory consists of $T=5$ time steps.
-In the quadrotor example presented in Section 4 of Musavi et al.
-(2024), the control input is specified as $u_{t}=\pi(x_{t})+\eta_{t}$, where $\pi(x_{t})$ denotes the nominal controller designed according to the approach in Alaimo et al.
-(2013).
-The system comprises $13$ states and $4$ control inputs.
-The unknown parameter matrix $\theta_{*}$ includes $7$ parameters, such as the mass $m$ and selected elements of the inertia matrix $I$.
-4.2 Estimation Error
+After discretization, the system dynamics can be rewritten in the structure of (1).
+In the quadrotor example presented in [10, Section 4], the control input is specified as $u_{t}=\pi(x_{t})+\eta_{t}$, where $\pi(x_{t})$ denotes the nominal controller designed according to the approach in [22].
+The system comprises 1313 states and $4$ control inputs.
+The unknown parameter matrix $\theta^{*}$ includes 77 parameters to learn.
+IV-B Estimation Error
 For each client $C_{i}$, the normalized error between the client and the server parameters is calculated as:
@@
 The maximum estimation error reported for communication round $r$ is: $e_{r}=\max_{1\leq i\leq M}e_{r}^{(i)}$, which also serves as a heterogeneity metric that quantifies how far the global model is from the most divergent client.
-5 Results
-Figure 3: Estimation error versus the number of global iterations for the real-world nonlinear dynamical system of a pendulum using gradient descent(GD).
-Results illustrate the impact of varying the number of clients ($M$), the number of local samples per client ($N_{i}$), and the heterogeneity parameter ($\epsilon$), with each client performing $K_{i}=1$ local updates at alearning rate of $10^{-2}$.
-Subfigures: (a) $N_{i}=10$, $\epsilon=0.01$; (b) $M=10$, $\epsilon=0.01$; (c) $M=20$, $N_{i}=10$.
-Figure 4: Estimation error vs.
-global iterations for the nonlinear pendulum system using mini-batch SGD (batch size 1010).
-Results illustrate the impact of varying the number of clients ($M$), the number of local samples per client ($N_{i}$), and the heterogeneity parameter ($\epsilon$), with each client performing $K_{i}=2$ local updates at learning rate of $10^{-3}$.
-Sub-figures: (a) $N_{i}=10$, $\epsilon=0.01$; (b) $M=10$, $\epsilon=0.01$; (c) $M=10$, $N_{i}=10$.
-Figure 5: Estimation error on synthetic data as a function of global iterations using gradient descent, evaluated across different client configurations of the number of clients ($M$), local dataset size per client ($N_{i}$), and heterogeneity parameter ($\epsilon$).
-In all cases, each client performs $K_{i}=5$ local update steps with a fixed learning rate of $10^{-4}$.
-The following configurations are considered: (a) $N_{i}=10$, $\epsilon=0.1$ with varying $M$; (b) $M=25$, $\epsilon=0.1$ with varying $N_{i}$; (c) $M=25$, $N_{i}=25$ with varying $\epsilon$.
-Figure 3 demonstrates the experimental results for a real-world, nonlinear dynamical system a pendulum.
-Figure 3 demonstrates the substantial impact of client participation on estimation error.
-With a fixed number of trajectories per client ($N_{i}=10$) and low system heterogeneity ($\epsilon=0.01$), increasing the number of participating clients from 1 to 81 significantly reduces the estimation error.
-Figure 3 reveals the critical role of the data volume in convergence performance.
-When maintaining a fixed number of clients ($M_{i}=10$) and low system heterogeneity ($\epsilon=0.01$), increasing the number of trajectories per client accelerates error reduction.
-This scaling demonstrates how federated learning enables each client to benefit indirectly from the expanded data pool through collaborative learning.
-Figure 3 illustrates the adverse impact of system heterogeneity on convergence behavior.
-Under a fixed number of clients ($M_{i}=10$) and trajectories ($N_{i}=10$), increasing the heterogeneity parameter leads to higher estimation error.
-As $\epsilon$ grows, it introduces greater variability into the model, which in turn makes precise estimation more difficult and increases the overall error.
-Figure 4 demonstrates the system behavior under stochastic gradient descent implemented with a batch size of 10 and two local updates per client.
-This configuration modifies both the granularity of the local optimization and the number of iterations per aggregation round.
-The introduction of minibatch stochastic gradient descent (SGD), rather than full-batch updates, further injects stochasticity into the optimization trajectory and may accelerate convergence and improve generalization on unseen trajectories.
-Using two local training updates enhances the local update effect; this can accelerate the reduction in estimation error under heterogeneous conditions, but may also exacerbate client drift in more heterogeneous settings.
-We systematically analyze gradient descent and mini-batch SGD for federated nonlinear system identification, offering practical insights for choosing algorithms under communication and computational constraints.
-Results for the quadrotor system have been provided in the Appendix.
-Figure 5 presents empirical convergence results for federated nonlinear system identification conducted on synthetic datasets, where the state-input pairs are generated from a synthetic nonlinear dynamical system.
-Specifically, Figure 5 illustrates that increasing the number of clients enhances estimation accuracy, given a fixed number of trajectories per client $(N_{i}=25$) and a constant heterogeneity parameter $(\epsilon=0.1)$.
-Similarly, Figure 5 demonstrates that augmenting the number of trajectories per client improves estimation quality when both the number of clients $(M)$ and the heterogeneity parameter $(\epsilon=0.01)$ remain fixed.
-Conversely, Figure 5 indicates a degradation in estimation performance as the heterogeneity parameter increases, with the number of clients $(M)$ and trajectories per client $(N_{i})$ held constant.
-These observations collectively underscore the trade-offs between data distribution, quantity, and heterogeneity in federated nonlinear system identification.
-Figure 6 shows graphs depicting how the normalized estimation error ($e$) varies with the square root of the number of clients ($\sqrt{M}$) for different values of epsilon ($\epsilon$) and fixed number of trajectories per client $(N_{i})$ for the pendulum dynamics and synthetic data.
-Figure 6 adopts a configuration analogous to that in Figure 3, whereas Figure 6 employs the setup outlined in Figure 5.
-This empirical validation highlights how the non-asymptotic convergence rate in federated learning can be systematically improved and scaled by a factor of $\sqrt{M}$ in the low heterogeneity regime.
-Figure 6 illustrates that increasing the number of local updates per communication round can reduce communication overhead by lowering the frequency of global aggregations.
-However, beyond a certain point, the performance gains diminish as excessive local training in each global iteration introduces significant global model divergence across clients.
-The experimental setup for Figure 6 is similar to that of Figure 4, except for a variation in the number of local updates.
 Figure 4: Comparison of estimation error versus $\sqrt{M}$ on (a) a real-world pendulum system and (b) synthetic data.
@@
 (c) Impact of local updates $(K_{i})$ on estimation error.
-6 Conclusion
-We introduced a federated learning framework for identifying linearly parameterized nonlinear dynamical systems, with a particular focus on piecewise affine (PWA) models.
+V Results
+Figure 2 demonstrates the experimental results the pendulum.
+In Figure 2, as the number of clients increases, the estimation error decays faster.
+Figure 2 shows that data volumme per client per round slighly improves the convergence rate.
+Figure 2 illustrates the adverse impact of system heterogeneity on convergence behavior.
+Under a fixed number of clients ($M_{i}=10$) and trajectories ($N_{i}=10$), increasing the heterogeneity parameter the estimation error.
+Additional results for the quadrotor system can be found in the Appendix.
+Figure 3 presents empirical convergence results for on synthetic datasets.
+Figure 3 illustrates that increasing the number of clients enhances estimation accuracy, Figure 3 demonstrates that increasing the number of trajectories per client improves estimation quality.
+Conversely, Figure 3 indicates a degradation in estimation performance as the heterogeneity parameter increases.
+Figure 4 shows how the normalized estimation error ($e$) varies with the square root of the number of clients ($\sqrt{M}$) for the pendulum dynamics and synthetic data.
+This result validates that the non-asymptotic convergence rate proposed is empirically consistent in the low heterogeneity regime.
+Figure 4 illustrates that increasing the number of local updates per communication round can reduce communication overhead by lowering the frequency of global aggregations.
+However, beyond a certain point, the performance gains diminish as excessive local training in each global iteration introduces significant global model divergence across clients.
+VI Conclusion
+We introduced a federated learning framework for identifying linearly-parameterized nonlinear dynamical systems, with a particular focus on piecewise affine (PWA) models.
 Our theoretical analysis shows that the convergence error decreases as $1/\sqrt{M}$, where $M$ is the number of clients, thus having significant improvement in convergence as more clients collaborate.
@@
 Adaptive batch size strategies could also balance the trade-off between gradient noise and convergence speed.
-A key fundamental open problem is where $\phi$ is not known a priori.
+A key fundamental open problem is when $\phi$ is not known a priori.
 One could explore scenarios where $\phi$ itself is learned collaboratively alongside the model, forming an end-to-end variant of our proposed framework.
-Finally, while this work focused on PWA models, a natural extension is to consider linearly parameterized nonlinear systems via Koopman theory (Brunton et al., 2022), which is based on lifting the states into an infinite-dimensional feature space, where the dynamics evolve linearly.
+Finally, while this work focused on PWA models, a natural extension is to consider linearly-parameterized nonlinear systems via Koopman theory [23], which is based on lifting the states into an infinite-dimensional feature space, where the dynamics evolve linearly.
 While exact infinite-dimensional embeddings are impractical, one can explore learning a finite set of invariant eigenfunctions to approximate this transformation.
-Appendix A Appendix
-In this section, we provide definitions for the $\mathcal{K}$ and $\mathcal{KL}$ functions, proofs for Lemma 1, Proposition 1, Proposition 2, and Theorem 1.
-In the end, we provide the results for quadrotor experiments.
-Definition 3 (Classes $\mathcal{K}$, and $\mathcal{KL}$).
-The following function class definitions are cited from Kellett (2014).
-  1. 1.
-    A function $\alpha:[0,a)\to[0,\infty)$ (with $a\in(0,\infty]$) is of class $\mathcal{K}$ if it is continuous, strictly increasing, and $\alpha(0)=0$.
-  2. 2.
-    A function $\beta:[0,a)\times[0,\infty)\to[0,\infty)$ is of class $\mathcal{KL}$ if, for each fixed $t\geq 0$, the map $r\mapsto\beta(r,t)$ belongs to $\mathcal{K}$, and for each fixed $r\in[0,a)$, the map $t\mapsto\beta(r,t)$ is continuous, nonincreasing, and $\lim_{t\to\infty}\beta(r,t)=0$.
-A.1 Proof of Lemma 1
-Proof.
-See Theorem 1 in Musavi et al. (2024). This lemma simply changes notation to fit the federated setting. ∎
-A.2 Proof of Proposition 1
-Proof.
-Fix a client $i$ and a unit vector $v\in\mathbb{S}^{n_{\varphi}-1}$, and let $Y_{t}:=v^{\top}\varphi(x^{(i)}_{t},u^{(i)}_{t})$ for $t=0,\ldots,T_{i}-1$.
-By Lemma 1 the sequence $\{Y_{t}\}$ satisfies the $(1,s_{\varphi},p_{\varphi})$ block–martingale small-ball condition, so Proposition 2.5 of Simchowitz et al.
-(2018), with $k=1$ and $\alpha=1$, gives
-$\Pr\!\bigl{[}\textstyle\sum_{t}Y_{t}^{2}\leq s_{\varphi}^{2}\!\bigl{(}N_{i}p_{\varphi}-2\log(1/\delta)\bigr{)}\bigr{]}\,\leq\,\delta.$
-Take a $1/4$-net $\mathcal{N}$ of the unit sphere and apply the same bound to each $v\in\mathcal{N}$ with failure level $\delta/|\mathcal{N}|:=\delta$; a union bound gives that, with probability $\geq 1-\delta$, the inequality holds for every $v\in\mathcal{N}$.
-Because any unit $u$ is within $1/4$ of some $v\in\mathcal{N}$
-$\lambda_{\min}\bigl{(}\Phi^{(i)}\Phi^{(i)\!\top}\bigr{)}\;\geq\;\frac{1}{2}\,s_{\varphi}^{2}\bigl{(}N_{i}Tp_{\varphi}-2\log(9^{n_{\varphi}}/\delta)\bigr{)}.$
-If $N_{i}T\geq\tfrac{4}{p_{\varphi}}\bigl{[}n_{\varphi}\log 9+\log(1/\delta)\bigr{]}$, the component inside the bracket is at least $\tfrac{1}{2}N_{i}Tp_{\varphi}$, and replacing $p_{\varphi}$ by $1$ loses only another factor $2$, giving $\lambda_{\min}(\Phi^{(i)}\Phi^{(i)\!\top})\geq\tfrac{1}{2}s_{\varphi}^{2}N_{i}T$ with probability at least $1-\delta$.
-Finally, substituting $\delta/M$ for $\delta$ and union-bounding over the $M$ clients yields the solution.
-∎
-A.3 Proof of Proposition 2
-Proof.
-Conditioned on $\Phi^{(i)}$, the columns $w^{(i)}_{t}$ are independent, mean-zero, $\sigma_{w}$-sub-Gaussian vectors. For any $t=0,\dots,T-1$, the matrix-form Bernstein inequality gives
-$\Pr\!\Bigl{(}\bigl{\|}W^{(i)}(\Phi^{(i)})^{T}\bigr{\|}_{2}>t\Bigr{)}\;\leq\;2\exp\!\Bigl{(}-\frac{t^{2}}{2\sigma_{w}^{2}\,\|\Phi^{(i)}\|_{F}^{2}}\Bigr{)}.$ (13)
-Since $\|\phi(x,u)\|_{2}^{2}\leq b_{\phi}$, we have $\|\Phi^{(i)}\|_{F}^{2}\leq N_{i}T\,b_{\phi}\leq N_{i}T(n_{x}+n_{\phi})$. Choosing
-$t\;=\;4\sigma_{w}\sqrt{\,N_{i}T\bigl{(}n_{x}+n_{\phi}+\log(2M/\delta)\bigr{)}}$
-makes the right side of (13) less than or equal to $\delta/(2M)$. A union bound over $i\in[M]$ and some algebra completes the proof. ∎
-A.4 Proof of Theorem 1
-Proof.
-Consider the relation for client $i$
-$X^{(i)}_{+}\;=\;\theta^{(i)*}\,\Phi^{(i)}+W^{(i)}.$
-Subtracting $\bar{\theta}_{\mathrm{LSE}}\Phi^{(i)}$ and regrouping gives
-$X^{(i)}_{+}-\bar{\theta}_{\mathrm{LSE}}\Phi^{(i)}=(\theta^{(i)*}-\bar{\theta}_{\mathrm{LSE}})\,\Phi^{(i)}+W^{(i)}.$
-Right-multiplying by $(\Phi^{(i)})^{\top}$ and summing over $i$,
-$\sum_{i=1}^{M}(X^{(i)}_{+}-\bar{\theta}_{\mathrm{LSE}}\Phi^{(i)})(\Phi^{(i)})^{\top}=\sum_{i=1}^{M}(\theta^{(i)*}-\bar{\theta}_{\mathrm{LSE}})\,\Phi^{(i)}(\Phi^{(i)})^{\top}+\sum_{i=1}^{M}W^{(i)}(\Phi^{(i)})^{\top}.$
-Defining [^2] Because $\bar{\theta}_{\mathrm{LSE}}$ minimizes the pooled least-squares objective, the left side sum vanishes.
-$G:=\sum_{i}\Phi^{(i)}(\Phi^{(i)})^{\top}$ and $\bar{\theta}^{*}:=\tfrac{1}{M}\sum_{i}\theta^{(i)*},$ we obtain
-$G\,(\bar{\theta}_{\mathrm{LSE}}-\bar{\theta}^{*})^{\top}=-\sum_{i}(\theta^{(i)*}-\bar{\theta}^{*})\Phi^{(i)}(\Phi^{(i)})^{\top}-\sum_{i}W^{(i)}(\Phi^{(i)})^{\top}.$
-We can bound each term as follows:
-$G\succeq\tfrac{1}{2}\,s_{\phi}^{2}N_{\mathrm{tot}}\,I$
-by Proposition 1,
-$\sum_{i}^{M}W^{(i)}(\Phi^{(i)})^{T}\leq 4\sigma_{w}\sqrt{N_{i}T(n_{x}+n_{\phi}+log(2M/\delta))}$
-by Proposition 2, and
-[…]
-using Assumption 5  [^3] Notice that Assumption 5 implies that $\theta^{(i)*}-\bar{\theta}^{(i)*}||\leq\epsilon$, $\forall i\in[M]$. and that $\Phi^{(i)}(\Phi^{(i)})^{\top}||_{2}\leq N_{i}b_{\phi}$.
-By multiplying both sides by $G^{-1}$, applying the bounds, taking the norm of each side gives with probability greater than or equal to $1-3\delta$,
-$\bar{\theta}_{LSE}-\bar{\theta}^{*}||_{2}\leq\frac{8\sigma_{w}}{s_{\phi}^{2}}\sqrt{\frac{n_{x}+n_{\phi}+log(2M/\delta)}{\sum_{i=1}^{M}N_{i}T}}+\frac{b_{\phi}}{s_{\phi}^{2}}\epsilon.$
-Since Assumption 5 implies $\|\theta^{(i)*}-\bar{\theta}^{*}\|_{2}\leq\varepsilon/2$ for $i\in[M]$,
-$\|\bar{\theta}_{\mathrm{LSE}}-\theta^{(i)*}\|_{2}\leq\|\bar{\theta}_{\mathrm{LSE}}-\bar{\theta}^{*}\|_{2}+\varepsilon/2.$
-This yields the expected bound. ∎
-Figure 7: Estimation error versus the number of global iterations for the real-world nonlinear dynamical system of a quadrotor using gradient descent (GD).
-Results illustrate the impact of varying the number of clients ($M$), the number of local samples per client ($N_{i}$),with each client performing $K_{i}=5$ local updates at a learning rate of $10^{-1}$.
-Sub-figures: (a) $N_{i}=10$, $\epsilon=2$; (b) $M=5$, $\epsilon=5$.
-A.5 Results for quadrotor
-Another real-world system considered is the quadrotor, adapted from the model proposed in Musavi et al.
-(2024); Alaimo et al.
-(2013), with necessary modifications as described in the following.
-Let $p\in\mathbb{R}^{3}$ and $v\in\mathbb{R}^{3}$ denote the position and velocity of the quadrotor’s center of mass in the inertial frame, respectively.
-Let $\omega\in\mathbb{R}^{3}$ represent the angular velocity in the body-fixed frame, and $q\in\mathbb{R}^{4}$ represent the orientation of the quadrotor using a unit quaternion.
-The corresponding equations of motion for the quadrotor are given by:
-$\dfrac{d}{dt}\begin{pmatrix}p\\ v\\ q\\ \omega\end{pmatrix}=\begin{pmatrix}v\\ -ge_{z}+\frac{1}{m}Qf_{u}\\ \frac{1}{2}\Omega q\\ I^{-1}(\tau_{u}-\omega\times I\omega)\end{pmatrix}+w,$
-where $g$ is the gravitational constant, $m$ is the total mass of the quadrotor, $I=\emph{diag}(I_{xx},I_{yy},I_{zz})$ denotes the inertia matrix in the body-fixed frame, $f_{u}\in\mathbb{R}$ is the total thrust, $\tau_{u}\in\mathbb{R}^{3}$ is the total moment expressed in the body-fixed frame, and $e_{z}=(0,0,1)^{\top}$ is the unit vector along the inertial $z$-axis.
-There are seven unknown parameters in total, with heterogeneity introduced in a subset of these parameters.
-$\theta_{1}$ $=\frac{1}{m}+\gamma_{1}^{(i)},\quad\theta_{2}=\frac{1}{I_{xx}},\quad\theta_{3}=\frac{I_{yy}-I_{zz}}{I_{xx}},$
-$\theta_{4}$ $=\frac{1}{I_{yy}},\quad\theta_{5}=\frac{I_{zz}-I_{xx}}{I_{yy}},\quad\theta_{6}=\frac{1}{I_{zz}},$
-$\theta_{7}$ $=\frac{I_{xx}-I_{zz}}{I_{zz}}.$
-Here $\gamma_{1}^{(i)}\sim\mathrm{U}(0,\epsilon)$.
-Figure 7 demonstrates the experimental results for quadrotor.
-Figure 7(a) demonstrates the impact of client participation on estimation error.
-With a fixed trajectory count per client ($N_{i}=10$) and system heterogeneity ($\epsilon=2$), increasing the number of participating clients reduces the error.
-Figure 7(b) underscores the role of data volume in convergence.
-Keeping the number of clients fixed ($M_{i}=5$) and heterogeneity ($\epsilon=5$) as fixed, increasing trajectories per client significantly accelerates error reduction.
-The length of each trajectory is set to 10 timesteps.
-These results follow similar trends as seen in the experiments for pendulum and synthetic datasets presented in the main paper.
+A formal proof establishing that Algorithm 1 converges to the quantity characterized in Theorem 1 remains an important direction for future work.
+While empirical results demonstrate consistent convergence behavior, deriving rigorous theoretical guarantees is ongoing.
+In particular, future work will focus on establishing convergence guarantees for both GD and SGD settings, including sufficient conditions and convergence rates.
+Acknowledgment
+This work is supported in part by ANRF under PM ECRG Grant (ANRF/ECRG/2024/005912/ENS), in part by Wadhwani Impact Trust under Exploratory Research Grant (SB25261132EEWDWI009167), and in part by IIT Madras under New Faculty Initiation Grant (RF24251118EENFIG009167).
+The authors thank Nandan Sudarsanam for fruitful discussions.